# PartComposer: Learning and Composing Part-Level Concepts from Single-Image Examples

### v1

@@ 4. Experiments
In this section, we first introduce our experiment setup in terms of data and mask generation, implementation details, and comparison details. We then show the qualitative results and comparisons followed by quantitative evaluations.
@@ 4.1. Experiments Setup
We use both synthetic data and real-world images in our experiments. We generate the synthetic images using DeepFloyd IF (DeepFloydAI, 2023) and we collect real-world images mainly from renderings of 3D objects (RenderHub, [n. d.]).
@@
  • Automatic segmentation and labeling: We apply off-the-shelf segmentation models such as SAM (Kirillov et al., 2023) to produce over-segmented masks, followed by GPT-4o-based (Achiam et al., 2023) captioning and labeling to group and assign part identities.
  • Manual or direct annotation: We directly specify or provide part masks, either manually or from existing annotated assets (e.g., 3D renderings (RenderHub, [n. d.]) or synthetic data).
Our framework is compatible with any source of part-level supervision, including both automatic and manual pipelines. Since the main focus of our work is not on the mask generation itself, we treat this step as a pre-processing module and do not over-emphasis our effort on it.
@@
For all experiments, we use Stable Diffusion v2.1 (Rombach et al., 2022) as the pre-trained text-to-image diffusion model and apply LoRA (Hu et al., 2022) with rank 32 to the U-Net module. The concept predictor is a convolutional network that operates on denoised latents and outputs both classification and segmentation predictions. It consists of three convolution layers (with output channels 16, 32, and 64), followed by two parallel output heads: a classification head composed of two fully connected layers for predicting multi-label concept presence, and a segmentation head consisting of a $1\times 1$ convolution followed by bilinear upsampling to produce per-concept spatial masks. We follow a two-stage training scheme similar to Break-a-Scene (Avrahami et al., 2023), where we only update the token embeddings with a high learning rate ($10^{-4}$) in the first stage and fully update both text encoder and LoRA weights in the U-Net with a low learning rate ($10^{-6}$) in the second stage. More training and inference details are introduced in Appendix A.
@@
For finetune-based approaches in concept learning, we compare our methods with 3 representative methods: Break-a-Scene (Avrahami et al., 2023), which is representative for concept learning from a single image, PartCraft (Ng et al., 2024), which is representative for part-level concept learning from large datasets, and MuDI (Jang et al., 2024), which is representative for disentangling subject-level concepts. To align their setting with our task, we adapt the original Break-a-Scene input dataset into a single-image examples input manner and feed part-decomposed examples into PartCraft and MuDI pipeline. We also conduct comparison with PiT (Richardson et al., 2025), which train dedicated priors for different categories and encode and recompose part-level concepts at test time. We directly use their provided checkpoints and prepare corresponding examples to inference and compare with our results. More implementation details for the adaptation of their methods are explained in Appendix B.
@@ 4.2.1. Qualitative Evaluation
We provide qualitative evaluation and comparisons our pipeline with other methods. We use two tasks - intra-category and cross-category inputs, each with more than 5 distinctive examples, to demonstrate the effectiveness of our pipeline in part-level concept learning and mixing.
@@
We first demonstrate the part-level concept learning and composition capability of our model using inputs from same categories. The input subjects consist same part decompositions (e.g., we decompose a chair into 4 parts: armrest, seat back, legs, and seat in Figure 5.) Note that we do not require the part decompositions to be semantically meaningful since our method can handle any part-level visual concepts.
Figure 5. Comparison of concept composition results for 2 chairs using our approach, Break-a-Scene (BaS) (Avrahami et al., 2023), PartCraft (Ng et al., 2024), and MuDI (Jang et al., 2024). The input images and correspond part decompositions are shown in the left column. We illustrates 4 part composition across the two input images by selecting 2 parts from each image, which are shown on the top row. We show 2 random samples for each composition.
Figure 5 shows the qualitative comparison of our method with Break-a-Scene (Avrahami et al., 2023), PartCraft (Ng et al., 2024), and MuDI (Jang et al., 2024), where we aim to recompose concepts across two input images. We generate 4 samples for each part composition. Break-a-Scene (Avrahami et al., 2023) struggles to disentangle and recompose part-level concepts, resulting in mixed identity in different parts, presumably since it was designed to do subject-level concept learning. PartCraft struggles to learn and recompose part-level concepts from single-image examples, resulting in poor and inconsistent image generation quality, presumably since it was designed to train on a large dataset. MuDI struggles to learn effective part-level concepts and result in poor composition capability, presumably since it was designed for few subject-level concepts disentanglement. Our pipeline produce clear part-level concept disentanglement and clean composition capability.
To enable fair comparison with PiT (Richardson et al., 2025) on their pre-trained checkpoints, we conduct two experiments, one for in-distribution and one for out-of-distribution. We first use their training data for ”creature” category and compare our results with the results generated by PiT’s ”creature” checkpoint. We then use a chair example to compare our results with the results generated by their ”product” checkpoint, which is the closest related category in their provided checkpoints. Figure 6 illustrates the qualitative comparison results. Even for in-distribution examples (the ”creature” example which is directly from their training data), PiT may drastically alter the identity of some concepts while our method generally preserves decent concept identity. When using out-of-distribution data like chairs, their pipeline fail to encode and compose provided parts.
@@
Figure 8 illustrates the results for 2 different levels of part granularity. It demonstrates that our method is agnostic to the granularity of parts and can effectively learn a large amount of (more than the 4 parts per image) very fine-grained concepts from single-image examples. We also provide substantial additional results for over 10 different categories in Appendix C.1.
@@
To enable more creativity in part-level concepts composition, we evaluate our methods in learning and composing parts from objects in different categories, aiming to generate virtual objects. The first row in Figure 5 shows hybrid compositions from a wheelchair and a bed and the second row in Figure 5 shows hybrid compositions from a wheel chair and a chair. Our methods can generate creative virtual objects with different part compositions, preserving clear part-level identity with reasonable structural arrangements for most composition scenarios. More results are shown in Figure 1 and in appendix C.2.
@@ 4.2.2. Quantitative Evaluation
We quantitatively compare our pipeline with baselines in terms of concept preserving capability and overall image generation quality. We run experiments for 3 categories (chair, vehicle, and characters), each with 7 single-image examples, and take the averaged score. For a single-image example containing $m$ images and with $k$ part decomposition, we sample 36 images for each of the $m^{k}$ possible part combinations, resulting in a total of around $1$k image samples per example.
@@
Our goal is to evaluate whether the intended part-level concepts are preserved in the generated images. Several previous works (Avrahami et al., 2023; Jang et al., 2024) that focus on subject-level concept learning use detectors to locate the subject and then compare feature-wise distance (e.g., DINO (Caron et al., 2021)) on the located areas. However, automatically detecting fine-grained parts from subjects is hard and DINO feature comparison is not effective on parts that are small or thin. Thus, we follow the evaluation methods used in PiT (Richardson et al., 2025), which use a multi-modal LLM - QWen (Yang et al., 2025) to compare the generated images with intended parts and score the visual alignment. Given $k$ parts to compose a subject, we prompt QWen to score from 00 to $k$, indicating how many parts are preserved in a generated image. We then normalize the scores for all of the examples into a 0-5 scale and report in Table 4.2.2. Our method significantly outperforms baselines to preserve the most intended part-level concepts.
@@
Generation Quality. We evaluate the overall image quality for the generated images, using 2 metrics to evaluate both the overall samples quality and single image comparison quality. We use FID and KID score to evaluate the image quality by comparing the sampled images for all possible part combinations and reference image datasets for corresponding category. For KID, we randomly divide the samples into 20 batches. The reference images are generated using SDXL (Podell et al., 2023). We also use multi-model LLMs to evaluate the generated image quality by directly comparing the samples with the given example images. We use QWen to score the quality on a 1-5 scale on the 5 methods for each example, where the best quality method get score 5 and the worst quality method gets score 1. Table 2 shows the results. Our model generally achieves the best image quality when comparing to other models.
@@ 4.2.3. Flexibility and Scalability
We demonstrate the flexibility and scalability of out pipeline in 3 aspects: encoding background concepts together with part-level concepts, dealing with incomplete part combination prompts, and scaling to learn large number of concepts from more than 2 single-image examples. The left column in Figure 9 demonstrates the our pipeline can learn background concepts with part concepts and naturally place the composed object in the background. The right column in Figure 9 shows that our method can generate structural complete objects given incomplete part combinations and can have variety in unspecified parts. The last row in Figure 10 demonstrate that our method can directly learn large number of concepts (i.e., 16 concepts from 4 single-image examples). All previous works often struggle to learn over 5 concepts at the same time without using large datasets.
4.2.4. Ablation Study and Analysis
We conduct ablation study by removing the dynamic data synthesis method and the concept predictor. Figure 7 illustrates the ablation results. Disabling the dynamic data synthesis results in poor data utility and disabling the concept predictor results in concept disentanglement or missing. We also conduct visualization for the convergence of concept predictor and weight/scale experiments in Appendix D. We analyze the limitations of our method in Appendix E.
@@ 5. Conclusion
We have introduced PartComposer, a method for fine-grained concept learning that enables new visual generation capabilities through part-level combination and composition. We propose a mutual information maximization framework that guides text-to-image diffusion model customization pipelines to both disentangle input concepts and reduce ‘missing concept’ failure modes. Combined with a dynamic data synthesis procedure, PartComposer is able to learn how to disentangle and compose fine-grained part-level concepts from single-image examples. We found that our approach was able to produce natural part compositions, that preserved concept identity, given single-image examples, from both matching and distinct categories. One promising direction for future work would be to convert our 2D productions into fully realized 3D models, using inverse rendering or image-to-3D pipelines. Looking forward, we believe our method can benefit creative design and imagination workflows, producing novel ideations from easy to source inputs, that might help spark new inspiration and creations.
@@ Appendix A Training Details
We provide the detailed training configuration used in our experiments for learning part-level visual concepts from single-image examples. Our pipeline is trained using Stable Diffusion v2.1 (Rombach et al., 2022) as the base text-to-image model. We use a pair of chair images from Figure 3 as an illustrative example. Each chair is decomposed into four semantic parts: armrest, backrest, legs, and seat. These parts are tokenized into 8 learnable placeholder tokens, denoted as <v1> through <v8>, with each token corresponding to a specific part instance from the two training images.
@@
We want to note that the second stage training step number is just a rough reference which generally ensures that all concepts are well-learned, and removes background contents. In general, after 18,000 steps, the concept learning and re-composition results are already good enough.
@@
We perform inference using standard DDIM (Song et al., 2020) sampling with a pretrained Stable Diffusion v2.1 backbone, the optimized text encoder, and the learned LoRA weights. Given a compositional prompt (e.g., "A photo of a partial chair with <v2>, <v5>, <v7>, <v8>, on a clean white background."), the model decodes a final image using 50 DDIM steps. We use a commonly used guidance scale 7.5.
@@
Our pipeline also supports learning extract background concepts if user wants to place the re-composed object on specific background images. We use <bgX> to represent the background concepts and they are learned together with the part-level concepts. In our data dynamic data synthesis stage, we replace the white background with the given background images and keep all other operations as the same. We use the background loss $\lambda_{\text{bg}}$ with weight $\lambda_{\text{bg}}=0.01$ to train the background concepts. Figure 9 show the results for incorporating an indoor background with re-composed chairs. Our method can naturally blend the newly composed chair into the given background.

### v2

@@ 4. Experiments
In this section, we first introduce our experiment setup in terms of data and mask generation, implementation details, and comparison details. We then show the qualitative results and comparisons, followed by quantitative evaluations.
@@ 4.1. Experiments Setup
We use both synthetic data and realistic images in our experiments. We generate the synthetic images using DeepFloyd IF (DeepFloydAI, 2023), and we collect realistic images mainly from renderings of 3D objects (RenderHub, 2025).
@@
  • Automatic segmentation and labeling: We apply off-the-shelf segmentation models such as SAM (Kirillov et al., 2023) to firstly produce over-segmented masks, followed by GPT-4o-based (Achiam et al., 2023) captioning and labeling to group and assign part identities.
  • Manual or direct annotation: We directly specify or provide part masks, either manually segmenting the parts or from existing annotated assets.
Our framework is compatible with any source of part-level supervision, including automatic and manual pipelines. Since the main focus of our work is not on the mask generation itself, we treat it as a pre-process step and do not overemphasize our effort on it.
@@
We use Stable Diffusion v2.1 (Rombach et al., 2022) as the pre-trained text-to-image diffusion model and apply LoRA (Hu et al., 2022) with rank 32 to the U-Net module. The concept predictor is a CNN network that operates on denoised latents and outputs both classification and segmentation predictions. It consists of three convolution layers (with output channels 16, 32, and 64), followed by two parallel output heads: a classification head composed of two fully connected layers for predicting multi-label concept presence, and a segmentation head consisting of a $1\times 1$ convolution followed by bilinear upsampling to produce per-concept spatial masks. We follow a two-stage training scheme similar to Break-a-Scene (Avrahami et al., 2023), where we only update the token embeddings with a high learning rate ($10^{-4}$) in the first stage and fully update both text encoder and LoRA weights in the U-Net with a low learning rate ($10^{-6}$) in the second stage. More training and inference details are introduced in Supplemental A.
@@
For finetune-based approaches in concept learning, we compare our methods with 3 representative methods: Break-a-Scene (Avrahami et al., 2023), which is representative for concept learning from a single image, PartCraft (Ng et al., 2024), which is representative for part-level concept learning from large datasets, and MuDI (Jang et al., 2024), which is representative for disentangling subject-level concepts. To align their setting with our task, we adapt the original Break-a-Scene input dataset into a single-image examples input manner and feed part-decomposed examples into the PartCraft and MuDI pipeline. We also conduct comparisons with PiT (Richardson et al., 2025), which trains dedicated priors for different categories and encodes and recomposes part-level concepts at test time. We directly use their provided checkpoints and prepare corresponding examples to conduct inference and compare with our results. More details are explained in Supplemental B.
@@ 4.2.1. Qualitative Evaluation
We provide qualitative evaluation and comparisons of our pipeline with other methods. We use two tasks - intra-category and cross-category inputs, each with more than 5 distinctive examples, to demonstrate the effectiveness of our pipeline in part-level concept learning and composing.
@@
We first demonstrate the part-level concept learning and composition capability of our model using inputs from the same categories. The input subjects consist same part decompositions (e.g., we decompose a chair into 4 parts: armrest, seat back, legs, and seat in Fig. 5.) Note that we do not require the part decompositions to be semantically meaningful since our method can handle any part-level visual concepts.
Figure 5. Comparison of concept composition results for 2 chairs using our approach, Break-a-Scene (BaS) (Avrahami et al., 2023), PartCraft (Ng et al., 2024), and MuDI (Jang et al., 2024). The input images and corresponding part decompositions are shown in the left column. We illustrate 4 part compositions across the two input images by selecting 2 parts from each image, which are shown on the top row. We show 2 random samples for each composition.
Fig. 5 shows the qualitative comparison of our method with Break-a-Scene (Avrahami et al., 2023), PartCraft (Ng et al., 2024), and MuDI (Jang et al., 2024), where we aim to recompose concepts across two input images. We generate 4 samples for each part composition. Break-a-Scene (Avrahami et al., 2023) struggles to disentangle and recompose part-level concepts, resulting in mixed identity in different parts, presumably since it was designed to do subject-level concept learning. PartCraft struggles to learn and recompose part-level concepts from single-image examples, resulting in poor and inconsistent image generation quality, presumably since it was designed to train on a large dataset. MuDI struggles to learn effective part-level concepts and results in poor composition capability, presumably since it was designed for a few subject-level concepts disentanglement. Our pipeline produces clear part-level concept disentanglement and clean composition capability.
To enable fair comparison with PiT (Richardson et al., 2025) on their released checkpoints, we conduct two experiments, one for in-distribution and one for out-of-distribution. We first use their training data for the ”creature” category and compare our results with the results generated by PiT’s ”creature” checkpoint. We then use a chair example to compare our results with the results generated by their ”product” checkpoint, which is the closest related category in their provided checkpoints. Fig. 6 illustrates the qualitative comparison results. Even for in-distribution examples (the ”creature”, which is directly from their training data), PiT may drastically alter the identity of some concepts, while our method generally preserves decent concept identity. When using out-of-distribution data like chairs, their pipeline fails to encode and compose the provided parts.
@@
Fig. 10 illustrates the results for 2 different levels of part granularity. It demonstrates that our method is agnostic to the granularity of parts and can effectively learn a large amount of (more than 4 parts per image) very fine-grained concepts. We also provide substantial results for over 10 other categories in Supplemental C.1.
@@
To enable more creativity in part-level concepts composition, we evaluate our methods in learning and composing parts from objects in different categories, aiming to generate virtual objects. The first row in Fig. 5 shows hybrid compositions from a wheelchair and a bed, and the second row in Fig. 5 shows hybrid compositions from a wheelchair and a chair. Our methods can generate creative virtual objects with different part compositions, preserving clear part-level identity with reasonable structural arrangements for most composition scenarios. More results are shown in Fig. 1 and in Supplemental C.2.
@@ 4.2.2. Quantitative Evaluation
We quantitatively compare our pipeline with baselines in terms of concept preserving capability and overall image generation quality. We run experiments for 3 categories (chair, vehicle, and characters), each with 7 single-image examples, and take the averaged score. For a single-image example containing $m$ images and with $k$ part decomposition, we sample 36 images for each of the $m^{k}$ possible part combinations, resulting in a total of around 11k image samples per example.
@@
Our goal is to evaluate whether the intended part-level concepts are preserved in the generated images. Several previous works (Avrahami et al., 2023; Jang et al., 2024) that focus on subject-level concept learning use detectors to locate the subject and then compare feature-wise distance (e.g., DINO (Caron et al., 2021)) on the located areas. However, automatically detecting fine-grained parts from subjects is hard, and DINO feature comparison is not effective on parts that are small or thin. Thus, we follow the evaluation methods used in PiT (Richardson et al., 2025), which use a multi-modal LLM - QWen (Yang et al., 2025) to compare the generated images with intended parts and score the visual alignment. Given $k$ parts to compose a subject, we prompt QWen to score from 0 to $k$, indicating how many parts are preserved in a generated image. We then normalize the scores for all of the examples into a 0-5 scale and report in Table 1. Our method significantly outperforms baselines to preserve the most intended part-level concepts.
@@
Generation Quality.
We evaluate the overall image quality for the generated images, using 2 metrics to evaluate both the overall samples’ quality and single image comparison quality. We use FID and KID scores to evaluate the image quality by comparing the sampled images for all possible part combinations and reference image datasets for the corresponding category. For KID, we randomly divide the samples into 20 batches. The reference images are generated using SDXL (Podell et al., 2023). We also use multi-model LLMs to evaluate the generated image quality by directly comparing the samples with the given example images. We use QWen to score the quality on a 1-5 scale on the 5 methods for each example, where the best quality method gets a score of 5 and the worst quality method gets a score of 1. Table 2 shows the results. Our model generally achieves the best image quality when compared to other models.
Table 2. Generation quality comparison for 3 categories.
@@ 4.2.3. Flexibility and Scalability
We demonstrate the flexibility and scalability of our pipeline in 3 aspects: encoding additional background concepts, dealing with incomplete part combination prompts, and scaling to learn a large number of concepts from more than 2 single-image examples. The left column in Fig. 11 shows that our pipeline can learn additional background concepts and naturally blend the composed object in the background. The right column in Fig. 11 shows that our method can generate structurally complete objects given incomplete part combinations and can have variety in unspecified parts. The last row in Fig. 12 shows that our method can directly learn a large number of concepts (i.e., 16 concepts from 4 single-image examples). All previous works often struggle to learn over 5 concepts at the same time without using large datasets.
4.2.4. Ablation Study
We conduct ablation studies by removing the dynamic data synthesis method and the concept predictor. Fig. 7 illustrates the ablation results. Disabling the dynamic data synthesis results in poor data utility, and disabling the concept predictor results in concept disentanglement or missing.
@@ 4.2.4. Ablation Study
4.2.5. Weight and Scale Experiments
The most important weight in our experiments is the weight for the concept predictor loss $\lambda_{\text{info}}$ explained in Section 3. We found that $\lambda_{\text{info}}=0.05$ is usually the best weight for most examples, though some examples require slightly higher weights ranging from $0.075$ to $0.1$. Increasing the $\lambda_{\text{info}}$ might show obvious improvement on concept preserving but may result in unnatural artifacts since the diffusion prior may get damaged. Fig. 8 illustrates this phenomenon. When increasing the weight from $0.01$ to $0.05$, concepts identity preservation is improved, showing the effectiveness of our concept predictor. When further increasing the weight, although concepts are still well-preserved, there are some visual artifacts, including saturation shift.
Figure 8. Visualization for the impacts of different information loss weights.
4.2.6. Visualization for concept predictor
We visualize the convergence of our concept predictor’s segmentation head to demonstrate the effectiveness of maximizing mutual information. Fig. 9 illustrates the comparison of segmentation results from the concept predictor and the ground truth masks. The concept predictor gradually converges to predict the correct segmentation of the parts, showing the effectiveness of our concept predictor design.
Figure 9. Visualization for the convergence of the segmentation head in our concept predictor.
5. Limitations
Although our method can learn well-entangled part-level concepts and re-compose them, the image generation quality for some challenging inputs still need improvement. For example, when composing parts from a truck and a sports car in Supplemental Fig.5, there exist artifacts in the generated images even though all the intended part concepts are preserved. Our pipeline also sometimes does not preserve the exact details for very thin structures, such as the number of bird legs (in Supplemental Fig.9). This problem might be partially due to the inherent limitation of text-to-image diffusion models, since these models can produce images containing unrealistic details. Additionally, for cross-category part re-composition, not all compositions of parts result in meaningful virtual objects. For example, when mixing a chair and a gym equipment in Supplemental Fig.11, the composition in the third column results in non-meaningful objects. A composition prediction scheme may be developed in this case to predict the meaningful part compositions.
@@ 6. Conclusion
We have introduced PartComposer, a method for fine-grained concept learning that enables new visual generation capabilities through part-level combination and composition. We propose a mutual information maximization framework to enable clear disentanglement of concepts and reduce ‘missing concept’ failure modes. Combined with a dynamic data synthesis procedure, PartComposer is able to learn how to disentangle and compose part-level concepts from single-image examples. Our approach is able to produce natural part compositions which preserve concept identity, from both matching and distinct categories. One promising direction for future work would be to convert our 2D productions into fully realized 3D models, using inverse rendering or image-to-3D pipelines. Looking forward, we believe our method can benefit creative design and imagination workflows, producing novel ideations from easy-to-source inputs, that might help spark new inspiration and creations.
Acknowledgements.
We would like to thank all the anonymous reviewers for their helpful suggestions. We would like to thank Caitlin Gong and Clara Fee for generating several single-image examples with masks and annotations. We would like to thank all the people whom we had discussions with for thinking about and commenting on interesting and illustrative examples to showcase the part re-composition capability. This work was funded by NSF award #1941808. Daniel Ritchie is an advisor to Geopipe and owns equity in the company. Geopipe is a start-up that is developing 3D technology to build immersive virtual copies of the real world with applications in various fields, including games and architecture.
@@ Appendix A Training Details
We provide the detailed training configuration used in our experiments for learning part-level visual concepts from single-image examples. Our pipeline is trained using Stable Diffusion v2.1 (Rombach et al., 2022) as the base text-to-image model. We use a pair of chair images from Fig. 3 in the main paper as an illustrative example. Each chair is decomposed into four semantic parts: armrest, backrest, legs, and seat. These parts are tokenized into 8 learnable placeholder tokens, denoted as <v1> through <v8>, with each token corresponding to a specific part instance from the two training images.
@@
We want to note that the second stage training step number is just a rough reference, which generally ensures that all concepts are well-learned, and removes background contents. In general, after 18,000 steps, the concept learning and re-composition results are already good enough.
@@
We perform inference using standard DDIM (20_DDIM) sampling with a pretrained Stable Diffusion v2.1 backbone, the optimized text encoder, and the learned LoRA weights. Given a compositional prompt (e.g., "A photo of a partial chair with <v2>, <v5>, <v7>, <v8>, on a clean white background."), the model decodes a final image using 50 DDIM steps. We use a commonly used guidance scale 7.5.
@@
Our pipeline also supports learning to extract background concepts if the user wants to place the re-composed object on specific background images. We use <bgX> to represent the background concepts, and they are learned together with the part-level concepts. In our data dynamic data synthesis stage, we replace the white background with the given background images and keep all other operations as the same. We use the background loss $\lambda_{\text{bg}}$ with weight $\lambda_{\text{bg}}=0.01$ to train the background concepts. Fig. 11 in the main paper shows the results for incorporating an indoor background with re-composed chairs. Our method can naturally blend the newly composed chair into the given background.